government claims that the stubble burning in the near by states follows a poisson process with a mean of 20 per day a) Estimate the probability of no burning per day


1
Expert's answer
2021-12-07T13:42:40-0500

Given,

Stubble burning per day "(\lambda)= 20"

Probability of no burning per day ?

"K=0"

So, the required probability"P(X=0)=\frac{e^{-\lambda}\times \lambda^k}{k!}"

Now, substituting the values,

"P(X=0)=\frac{e^{-20}\times 20^0}{0!}\\ \\
\Rightarrow P(X=0)=\frac{e^{-20}}{1} =e^{-20}"